OpenStudy (anonymous):

Suppose x coins are tossed. Write an expression to represent the number of possible outcomes

OpenStudy (anonymous):

When \(x=1\) you have two outcomes. As \(x\) increases by \(1\), the outcomes double.

OpenStudy (anonymous):

i know wio it's a stupid qestion

OpenStudy (anonymous):

This is a geometric progression.

OpenStudy (anonymous):

You asked it!

OpenStudy (anonymous):

no it's one of my qestion's

OpenStudy (anonymous):

If you know the answer, you don't need to come here.

OpenStudy (anonymous):

i don't know that's why im here

OpenStudy (anonymous):

Well\[ a_n = a_1 r^{n-1} \]We know \(r=2\) and \(a_1=2\)

OpenStudy (anonymous):

\[ a_n = (2)(2)^{n-1}=2^n \]

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

Should be using \(x\) instead of \(n\).... \[ \text{# outcomes} = 2^x \]

OpenStudy (anonymous):

thnkz a lot best responce me

OpenStudy (anonymous):

....... Why? You didn't even give me best response!

OpenStudy (anonymous):

i did gime on plz

OpenStudy (anonymous):

.....